# Data Caricatures: On the Representation of African American Language in Pretraining Corpora

### v1

@@ 5.1 Methods
Human Judgments. For each text that is judged to be human written and contain identifiable AAL features, annotators are asked to also rate each text on three additional dimensions. First, we follow Deas et al. (2024) and include Native Speaker which asks whether the text appears to be naturally written by an AAL-speaker. The remaining two dimensions capture whether the AAL texts are suitable for training LLMs. Inspired by Fleisig et al. (2024), Stereotype asks whether the text portrays a stereotypical or harmful representation of AAL or its speakers. Finally, we consider annotators’ own linguistic attitudes toward texts that are included in pretraining corpora, motivated by sociolinguistics work on listeners’ language attitudes and perceptions Campbell-Kibler (2008). We include Appropriateness which asks whether an annotator perceives a given text to be appropriate for language models to generate and thus appropriate to include in pretraining data.
All dimensions are again judged on 4-point Likert scales with higher values representing that a text is judged to be likely written by a native speaker, to perpetuate stereotypes, and to be appropriate for an LLM to generate. Annotators exhibit substantial agreement for Native Speaker ($\kappa$ = 0.619) although annotators’ varying personal perspectives on the Appropriateness and Stereotype dimensions expectedly yield little to no agreement ($\kappa=$ 0.188 and -0.021 respectively), highlighting the need for representation of AAL-speakers in model development processes. Additional details and discussion of agreement are included in Appendix D.
Misrepresentative Language. To explore whether the representation of AAL in C4 is indicative of naturalistic AAL, we first estimate the presence of hip hop and rap lyrics in C4 variants. We particularly focus on song lyrics because large-scale corpora of published hip hop lyrics are available and enable measuring overlap with each pretraining corpus. Leveraging the aforementioned human judgments, we also estimate the perceived presence of AAL use by non-AAL-speakers.
We follow the deduplication approach of Brown et al. 2020 which measures the overlap of 8-13 token n-grams. While an ideal analysis would measure exact matches, through manual examination, we find instances of hip hop lyrics embedded in non-lyric text, modifications to lyrics such as censored terms, and orthographic differences (e.g., "going" vs. "goin") that would not be captured by such an approach. We note that while we use a large corpus of song lyrics, a comprehensive dataset remains infeasible and our estimates likely reflect lower bounds. Methodological details are included in Appendix F and additional analyses are included in Appendix G.
@@ 5.2 Results
Figure 3: % of AAL documents extracted from C4 containing shared n-grams with hip hop lyrics for varying n-gram token lengths. Tokenization uses the Llama-2 tokenizer.
@@
Hip Hop Lyrics. Figure 3 presents the percentage of documents overlapping with hip hop lyrics by n-gram token length. Nearly 15% of C4.en.noBlocklist and 12% of C4.en documents contain substantial overlap (8-grams) with hip hop lyrics. Examples of lyrics identified by annotators are included in Table 3. While there are AAL features present, such as the use of negative concord in None of my cars ain’t rented, any inclusion of hip hop lyrics may be misrepresentative of how AAL-speakers would interact with LLMs.
Non-AAL-Speakers. AAL may also be misrepresented through use by non-AAL-speakers, predominantly online. While it is infeasible to automatically determine the linguistic background of a text’s author, among the C4.en texts annotators judged to contain AAL features, 51% were judged to have been unlikely written by a native AAL-speaker. Most of these texts (70%) were perceived as not human-written.  [^6] As an approximate reference, only 21% of human-written AAL texts in Deas et al. (2023) were judged to be likely not human-written or unnatural. The remaining texts, however, include corporate social media accounts adapting AAL-like language (e.g., …this will get you where you need to be. Ball out by clicking the link below.) as well as online forum posts (e.g. Dat be that real deal Hip Hop! …Put some spect on it!). These texts often exaggerate the use of AAL features, contributing to misunderstandings of native speakers. We conclude that a significant portion of texts with AAL features are unlikely to have been written by an individual AAL-speaker, and misrepresent naturalistic AAL.
Many features of AAL are shared with other language varieties and the appearance of shared features in different linguistic contexts may complicate LLMs’ ability to model these varieties during pretraining. Annotators noted many of these cases, such as the Hawai’ian Creole text …An make us shame cuz we no mo husban. Dat Yahweh make come up from da king ohana… which may be misinterpreted as Digital AAL (DAAL; Cunningham (2014)). [^7] See Appendix H for further discussion.
Figure 4: Proportion of extracted AAL from C4 judged to be inappropriate for models to generate and judged to reinforce stereotypes of AAL or its speakers. Proportions are show separately for the filtered and unfiltered subsets of the extracted data.
@@
Innapropriateness and Stereotyping. The results of Appropriateness and Stereotype judgments are shown in Figure 4. First, we find that the filtering applied to C4.en.noBlocklist removes many inappropriate and stereotype-reinforcing texts as intended, with C4.en.noBlocklist exceeding the rates of C4.en. We also find, however, that substantial proportions of remaining texts in C4.en are judged to be inappropriate and stereotype-reinforcing–28.9% and 24.5% respectively. Many of these texts contain variants of C4 block list terms or terms self-censored with asterisks. For example, the first two texts in Table 4 are judged to be inappropriate for models to generate and contain references to block list terms (****in and fuk).
@@ 6 RQ3: How does filtering impact AAL representation?
The previous experiments show that AAL is often underrepresented and misrepresented in modern, open-source pretraining corpora. We follow these experiments with an evaluation of automated filtering strategies used in pretraining corpora to better understand potential underlying determinants of AAL representation.
@@ 6.1 Methods
Table 5: Summary of datasets used in filtering experiments. We sample equal numbers of texts to represent AAL and WME, and where possible, sample texts of similar length.
The C4 block list is known to disproportionately filter AAL from pretraining data Dodge et al. (2021). To investigate how recent, model-based language, toxicity, and quality filters affect AAL representation, we evaluate 16 automated filters on AAL texts. A full list with descriptions of each filter evaluated is included in Appendix I.
@@
RedPajama Filtering RedPajama-v2 Weber et al. (2024) is a large-scale pool of Common Crawl text with pre-computed filtering heuristics. It was introduced to enable efficient filtering rather than to directly pretrain LLMs without further pre-processing. To evaluate automated filters, we extract all texts with an African American-aligned English probability $\geq 0.8$ and randomly sample an equal number of texts with a White-aligned English probability of $\geq 0.8$ following prior work Xia et al. (2020).
@@ 6.2 Results
Figure 5: Average normed filter probabilities given by each automated filter for the positive label across AAL and WME in RedPajama (top) and across data sources (bottom). Error bars represent 99% confidence intervals, and (top) indicates significant differences ($p<0.01$). Shaded columns indicate filters that prefer AAL texts (top) and prefer Hip Hop/AAL Tweets (bottom). Larger values indicate texts are more likely to be preserved.
@@
AAL Sources. To better understand how automated filters’ predictions vary across AAL sources, Figure 5 (bottom) presents the same average normalized model predictions on AAL texts from dialogues (CORAAL), song lyrics (Hip Hop), and social media (AAL Tweets). As CORAAL contains natural speech and is unlikely to contain the lexical features exhibited by hip hop lyrics and social media, most filters (11 of 16 ) expectedly are more likely to conserve dialogue transcripts over the other sources. Most filters (12 of 16 ) are also more likely to conserve AAL tweets over hip hop lyrics. While filters appear to elevate natural speech, texts like transcripts in CORAAL are not widely available online suggesting that the sources of pretraining data may be the more influential factor determining AAL’s representation in corpora (see Appendix J for further discussion).
@@ 7 Discussion and Conclusion
In this work, we examine the quantity and quality of AAL representation in pretraining corpora as well as the impact of recent automated quality filters. We show that AAL is widely underrepresented in modern corpora, and its representation is highly sensitive to choices made during dataset construction–namely data sources and filtering. Underlying these findings, language understanding benchmarks that drive LLM development are unlikely to contain AAL, potentially incentivizing the similarly minimal AAL representation found in pretraining corpora. Poor representation may reinforce the discriminatory language model behaviors found in prior work Fleisig et al. (2024); Hofmann et al. (2024) and restrict the benefits of LLMs for AAL-speakers Brewer et al. (2023).
Driven by our findings and prior work, incorporating understanding of text sources in LLM development and carefully curating these sources through inclusion of AAL-speakers as stakeholders in design processes Friedman (1996); Suresh et al. (2024) is necessary to faithfully represent naturalistic AAL use. We recommend that in developing LLMs designed for socially impactful domains where AAL-speakers already face disparate impacts (e.g., medical Beach et al. (2021) or legal Jones et al. (2019) settings), pretraining data curators ensure that the sources and filtering processes retain representative AAL data. Such precautions may also accordingly be extended to the benchmarks used to evaluate the effectiveness of design decisions throughout the LLM development cycle, incentivizing understanding of sociolinguistic variation as well as commonsense understanding and other capabilities.
@@ 9 Ethics Statement
As a significant portion of this work involves qualitative analyses of texts, estimates, we acknowledge that the authors’ backgrounds and perspectives may impact interpretation of some findings. In this work, we attempt to calculate and report quantitative statistics to complement qualitative discussion of pretraining corpora’s contents. As we strictly analyze open-sourced pretraining corpora, we encourage readers to examine texts within each corpus as well as form their own interpretations from the presented results.
@@
Our findings contribute to the ongoing study of AAL biases in both language understanding and generation systems. Prior work highlights the barriers that language technologies’ lack of AAL understanding poses to African American users Cunningham et al. (2024), while others detail how human prejudices can lead to, for example, disproportionately labeling AAL as toxic Sap et al. (2022). With respect to language generation, a recent study has found that African American users prefer chatbots that use WME rather than AAL in generations Finch et al. (2025). These findings highlight the conflict between language technologies that are able to properly understand AAL but not necessarily to generate AAL. As such, while our findings document factors in pretraining data that may lead to misunderstanding of AAL in LLMs, increasing representation alone does not absolve language model creators of their obligation to take in the culturally relevant context.
To conclude, no LLM depiction of AAL should misrepresent and caricature AAL use. In contexts where LLM use serves the social benefit of AAL-speakers, measures should be taken to ensure that depictions of AAL do not trivialize the rich linguistic and cultural background surrounding them.
@@ Appendix A Large-Scale Corpora Analysis
We select 12 open-source pretraining corpora for our analyses, prioritizing those that have been used to pretrain models often used in academic research. Given our focus on AAL, we also focus analyses corpora that are intended for pretraining predominantly English rather than multilingual models. For all corpora, we strictly analyze portions that are publicly available and do not include subsets of RedPajama Computer (2023) or The Pile Gao et al. (2020) that are restricted due to copyright. All corpora are only used for academic research purposes in line with their intent.
@@ A.1 Pretraining Corpus Scanning
Our data pipeline is written in Python, making use of the Polars library Vink et al. (2025) for performance-sensitive data processing, as well as Pandas pandas development team (2020) and MatPlotlib Hunter (2007) for analyses and visualization. Labeling of independent subsets of the data is a trivially parallelizable task. We partition the corpus into disjoint splits and use 20 independent worker processes, one for each available physical CPU core to store the full text of records matching our selection criterion. In all, we obtain labels for over 16 TB of compressed records, as stored on disk. For capturing approximations of full distributions, we additionally use $t$-digest Davidson-Pilon (2015).
@@ A.2 Corpus Sample Analyses
Our file sampling policy was motivated by the constraints of the available metadata, the size of the corpora, and the goal of our analysis being reproducible. File sizes varied between the corpora and within corpora, so a fixed number of files would create an in. With the sheer amount of variance in the metadata of the corpora, the only piece of metadata common to all corpora was the URL hosting the data file. To create a uniform random sampling of the files in the corpora, we use a variant on hash-based consistent sampling Manasse et al. (2010). We hash the full url hosting the data with sha256, sample the first $n$ files sorted by hash greater than our threshold of 250 GB per corpus, due to constraints on our resources.

### v2

@@ 5 RQ2: How Well is AAL Represented?
After characterizing the extent to which AAL is represented in open-source pretraining corpora, in this section, we aim to characterize the quality of the included AAL documents. To do so, we use the AAL subsets extracted in the previous experiments to further analyze the sources, naturalness, and inoffensiveness of the AAL texts.
@@ 5.1 Methods
Human Judgments. For each text that is judged to be human written and contain identifiable AAL features, annotators are asked to also rate each text on three additional dimensions. First, we follow Deas et al. (2024) and include Native Speaker which asks whether the text appears to be naturally written by an AAL speaker. With the remaining two dimensions, we capture annotators’ perceptions and linguistic attitudes toward a given text considering its inclusion in pretraining data. Inspired by Fleisig et al. (2024), Stereotype first asks whether the text portrays a stereotypical or harmful representation of AAL or its speakers. Additionally, we include Appropriateness which asks whether an annotator perceives a given text to be appropriate for language models to generate, and therefore, appropriate to include in pretraining data. Because we are interested in capturing the annotators’ own perspectives on pretraining texts and their interpretations of these dimensions, we avoid providing detailed definitions of Stereotype and Appropriateness dimensions in particular, motivated by sociolinguistics work on language attitudes and perceptions Campbell-Kibler (2008); Labov et al. (2011).
All dimensions are again judged on 4-point Likert scales with higher values representing that a text is judged to likely have been written by a native speaker, to perpetuate stereotypes, and to be appropriate for an LLM to generate. Annotators exhibit substantial agreement for Native Speaker ($\kappa$ = 0.619) although annotators’ varying personal perspectives on the Appropriateness and Stereotype dimensions yield little to no agreement ($\kappa=$ 0.188 and -0.021 respectively). Given that these dimensions are highly dependent on individuals’ own perceptions, this level of agreement is expected. Additionally, these results highlight the need to ensure that the inclusion of AAL speakers’ perspectives captures the variety of perspectives. Additional details and discussion of agreement are included in Appendix D.
Misrepresentative Language. To explore whether the representation of AAL in C4 is indicative of naturalistic AAL, we first estimate the presence of language resembling hip hop and rap lyrics in C4 variants. We particularly focus on song lyrics because large-scale corpora of published hip hop lyrics are available, enabling the measurement of overlap with each pretraining corpus. Leveraging the aforementioned human judgments, we also estimate the perceived presence of AAL use by non-AAL speakers.
We follow the deduplication approach of Brown et al. 2020 which measures the overlap of 8-13 token n-grams. While an ideal analysis would measure exact matches, through manual examination, we find instances of hip hop lyrics embedded in non-lyric text, censored terms, and orthographic differences (e.g., "going" vs. "goin") that would not be captured by such an approach. We note that while we use a large corpus of song lyrics, a comprehensive dataset remains infeasible and our estimates likely reflect lower bounds. Methodological details are included in Appendix H.
@@ 5.2 Results
Figure 4: Percent of AAL documents extracted from C4 containing shared n-grams with hip hop lyrics for varying n-gram token lengths. Tokenization uses the Llama-2 tokenizer.
@@
Hip Hop Lyrics. Figure 4 presents the percentage of documents overlapping with hip hop lyrics by n-gram token length. Nearly 15% of C4.en.noBlocklist and 12% of C4.en documents overlap with hip hop lyrics (8-grams). Examples of lyrics identified by annotators are included in Table 3. While there are AAL features present, such as the use of negative concord in None of my cars ain’t rented, any inclusion of hip hop lyrics may be misrepresentative of how AAL speakers would interact with LLMs. Analyses of the remaining corpora are included in Appendix I, and additional related analyses are included in Appendix J.
Non-AAL-Speakers. AAL may also be misrepresented through use by non-AAL speakers, predominantly online. While it is infeasible to automatically determine the linguistic background of a text’s author, among the C4.en texts annotators judged to contain AAL features, 51% were judged unlikely to have been written by a native AAL speaker. Most of these texts (70%) were perceived as not human-written.  [^7] As an approximate reference, only 21% of human-written AAL texts in Deas et al. (2023) were judged to be likely not human-written or unnatural. The remaining texts, however, include corporate social media accounts adapting AAL-like language (e.g., …this will get you where you need to be. Ball out by clicking the link below.) as well as online forum posts (e.g. Dat be that real deal Hip Hop! …Put some spect on it!). These texts often exaggerate the use of AAL features, contributing to misunderstandings of AAL speakers. We conclude that a substantial portion of texts with AAL features are unlikely to have been written by an individual AAL speaker, and misrepresent naturalistic AAL.
Many features of AAL are shared with other language varieties and the appearance of shared features in different linguistic contexts may complicate LLMs’ ability to model these varieties during pretraining. Annotators noted many of these cases, such as the Hawai’ian Creole text "…An make us shame cuz we no mo husban. Dat Yahweh make come up from da king ohana…" which may be misinterpreted as Digital AAL Cunningham (2014) [^8] See Appendix K for further discussion.
Figure 5: Distribution of annotator ratings on stereotype and appropriateness dimensions. Shades correspond to Likert scale values. Percentages shown in white reflect the proportion of texts judged to be inappropriate and to reinforce stereotypes (outlined in black).
@@
Innapropriateness and Stereotyping. Appropriateness and Stereotype judgment results are shown in Figure 5. First, we find that the filtering applied to C4.en.noBlocklist removes many inappropriate and stereotype-reinforcing texts as intended, with C4.en.noBlocklist exceeding the rates of C4.en. We also find, however, that substantial proportions of remaining texts in C4.en are judged to be inappropriate and stereotype-reinforcing–28.9% and 24.5% respectively. Many of these texts contain variants of C4 block list terms or terms self-censored with asterisks. For example, the first two texts in Table 4 are judged to be inappropriate for models to generate and contain references to block list terms (****in and fuk).
@@ 6 RQ3: How does filtering impact AAL representation?
The previous experiments show that AAL is often underrepresented and misrepresented in modern, open-source pretraining corpora. We follow these experiments with an evaluation of automated filters to better understand potential underlying determinants of AAL representation.
@@ 6.1 Methods
Table 5: Summary of filtering experiment datasets.
The C4 block list is known to disproportionately filter AAL from pretraining data Dodge et al. (2021). To investigate how recent, model-based language, toxicity, and quality filters affect AAL representation, we evaluate 16 automated filters on AAL texts. A full list with descriptions of each filter evaluated is included in Appendix L.
@@
RedPajama Filtering RedPajama-v2 Weber et al. (2024) is a large pool of Common Crawl text with pre-computed filtering heuristics. To evaluate automated filters, we extract all texts with an AAL posterior probability 0.8$\geq$ 0.8≥ 0.8 and randomly sample an equal number of texts with a White-aligned English probability of 0.8$\geq$ 0.8≥ 0.8 following prior work Xia et al. (2020).
Figure 6: Average outputs of automated filters between AAL and WME RedPajama texts (top) and across AAL data sources (bottom). Outputs are standardized with larger values indicating texts are more likely to be preserved in the corpus. Error bars represent 99% confidence intervals, and (top) indicates significant differences ($p<0.01$). Shaded columns indicate filters that prefer AAL texts (top) and prefer Hip Hop/AAL Tweets (bottom).
@@ 6.2 Results
AAL Sources. To better understand how automated filters’ predictions vary across AAL sources, Figure 6 (bottom) presents the same average normalized model predictions on AAL texts from dialogues (CORAAL), song lyrics (Hip Hop), and social media (AAL Tweets). As CORAAL contains natural speech and is unlikely to contain the lexical features exhibited by hip hop lyrics and social media, most filters (11 of 16) expectedly are more likely to conserve dialogue transcripts over the other sources. Most filters (12 of 16) are also more likely to conserve AAL tweets over hip hop lyrics. While filters appear to elevate natural speech, texts like transcripts in CORAAL are not widely available online suggesting that the sources of pretraining data may be the more influential factor determining AAL’s representation in corpora (see Appendix M for further discussion).
@@ 7 Discussion and Conclusion
In this work, we examine the quantity and quality of AAL representation in pretraining corpora as well as the impact of recent automated quality filters on such representation. We first show that AAL is widely underrepresented in modern corpora, and that many of those AAL documents are shared among corpora. We additionally highlight the ways in which AAL documents may be misrepresentative in these corpora, including analyses of highly performative language as well as documents that are judged to be written by non-AAL speakers, stereotypical, or inappropriate for LLMs’ to generate. Finally, we show how automated filters used in recent pretraining corpora impact the quantity and quality of AAL representation.
Our findings have several implications. The lack of AAL in pretraining corpora may underlie LLMs’ difficulties in understanding AAL (e.g., Groenwold et al. (2020); Deas et al. (2023)), particularly for specific features Kleiner et al. (2024); Grieser et al. (2024). Beyond its underrepresentation, the misrepresentation of AAL in pretraining corpora may reinforce the discriminatory language model behaviors that have been identified in prior work Fleisig et al. (2024); Hofmann et al. (2024) and restrict the benefits of LLMs for AAL speakers Brewer et al. (2023). Similar language technologies’ inability to reliably interpret AAL has also been shown to create barriers to use of the technology Cunningham et al. (2024); Harrington et al. (2022). Documenting the representation of AAL in pretraining corpora enables better understanding of why LLMs and other technologies pose such risks and offers insights into methods of mitigating them.
While we find that automated filters are more likely to remove AAL than WME, they are also more likely to conserve naturalistic speech (CORAAL) over tweets and hip hop lyrics. Considering we highlight through RQ2 that highly performative AAL is prevalent in the corpora, this also suggests that there is a lack of naturalistic AAL in the sources leveraged for pretraining corpora. Incorporating understanding of text sources in LLM development and carefully curating these sources through meaningful inclusion of AAL speakers as stakeholders in design processes Friedman (1996); Suresh et al. (2024) is necessary to faithfully represent naturalistic AAL use. We recommend that in developing LLMs designed for socially impactful domains where AAL speakers already face disparate impacts (e.g., medical Beach et al. (2021) or legal Jones et al. (2019) settings), pretraining data curators ensure that the sources and filtering processes produce representative AAL data. Finally, language understanding benchmarks that drive LLM development are unlikely to contain AAL, potentially leading to the similar underrepresentation of AAL found in pretraining corpora. The aforementioned considerations may also accordingly be extended to the benchmarks used to evaluate the effectiveness of design decisions throughout the LLM development cycle, incentivizing understanding of sociolinguistic variation as well as commonsense understanding and other capabilities.
@@ 9 Ethics Statement
As a significant portion of this work involves qualitative analyses of texts and estimates, we acknowledge that the authors’ backgrounds and perspectives may impact interpretation of some findings. In this work, we attempt to calculate and report quantitative statistics to complement qualitative discussion of the contents of pretraining corpora. As we strictly analyze open-source pretraining corpora, we encourage readers to examine texts within each corpus as well as form their own interpretations from the presented results.
@@
Our findings contribute to the ongoing study of AAL biases in both language understanding and generation systems. Prior work highlights the barriers that language technologies’ lack of AAL understanding poses to African American users Cunningham et al. (2024), while others detail how human prejudices can lead to, for example, disproportionately labeling AAL as toxic Sap et al. (2022). With respect to language generation, recent work has found that African American users prefer and trust current chatbots using WME rather than AAL in generations Finch et al. (2025); Basoah et al. (2025). These findings highlight the conflict between language technologies that are able to properly understand AAL but not necessarily to generate AAL. As such, while our findings document factors in pretraining data that may lead to misunderstanding of AAL in LLMs, increasing representation alone does not absolve language model creators of their obligation to take in the culturally relevant context.
To conclude, no LLM depiction of AAL should misrepresent or caricature AAL use. In contexts where LLM use serves the social benefit of AAL speakers, measures should be taken to ensure that depictions of AAL do not trivialize the rich linguistic and cultural background surrounding them.
10 Acknowledgments
This work was supported in part by a Provost Diversity Fellow mini-grant, a grant from the Knight First Ammendment Institute at Columbia University, and grant IIS-2106666 from the National Science Foundation. The first author is supported by the National Science Foundation Graduate Research Fellowship DGE-2036197, the Columbia University Provost Diversity Fellowship, the Columbia School of Engineering and Applied Sciences Presidential Fellowship. Any opinion, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation or Knight First Amendment Institute. We thank the anonymous reviewers and the following people for providing valuable feedback on an earlier draft: Elsbeth Turcan, Matthew Toles, and Narutatsu Ri.
@@ Appendix A Large-Scale Corpora Analysis
We select 12 open-source pretraining corpora for our analyses, prioritizing those that have been used to pretrain models often used in academic research. Given our focus on AAL, we also focus analyses on corpora that are intended for pretraining predominantly English rather than multilingual models. For all corpora, we strictly analyze portions that are publicly available and do not include subsets of RedPajama Computer (2023) or The Pile Gao et al. (2020) that are restricted due to copyright. All corpora are only used for academic research purposes in line with their intent.
@@ A.1 Pretraining Corpus Scanning
Our data pipeline is written in Python with Polars Vink et al. (2025) for performance-sensitive data processing, Pandas pandas development team (2020), UpSetPlot Lex et al. (2014), and MatPlotlib Hunter (2007) for analyses and visualization. Labeling of independent subsets of the data is a trivially parallelizable task. We partition the corpus into disjoint splits and use 20 independent worker processes, one for each available physical CPU core to store the full text of records matching our selection criterion. In all, we obtain labels for over 16 TB of compressed records, as stored on disk. For capturing approximations of full distributions such as in 8, we additionally use $t$-digest Davidson-Pilon (2015).
@@ A.2 Corpus Sample Analyses
Our file sampling policy was motivated by the constraints of the available metadata, the size of the corpora, and the goal of our analysis being reproducible. File sizes vary between corpora and within corpora, so a fixed number of files would not result in balanced representation. Due to the low metadata consistency of the corpora, the only common attribute to all corpora was the URL hosting the data file. To create a uniform random sampling of the files in the corpora, we use a variant on hash-based consistent sampling Manasse et al. (2010). Thus, the selection procedure entailed hashing the full url hosting the data with sha256, then sampling the first $n$ files (sorted by hash) until a threshold of 250 GB was reached for each corpus.